OpenStudy (anonymous):

plz help check my answer! Pyramid A is a square pyramid with a base side length of 9 inches and a height of 10 inches. Pyramid B has a volume of 1350 cubic inches. How many times bigger is the volume of pyramid B than pyramid A? Give your answer as a percent. Provide an explanation and proof for your answer to receive full credit. is it 45%

OpenStudy (anonymous):

is it 45%

OpenStudy (anonymous):

@phi @ganeshie8 @mathstudent55 @jdoe0001 @thomaster

OpenStudy (anonymous):

@zzr0ck3r

jimthompson5910 (jim_thompson5910):

What is the volume of pyramid A? What is the volume of pyramid B?

OpenStudy (anonymous):

pyramid b is shown and a is 30 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

B is shown?

OpenStudy (anonymous):

yes its 1350

jimthompson5910 (jim_thompson5910):

pyramid A is incorrect

OpenStudy (anonymous):

why? can u explain

jimthompson5910 (jim_thompson5910):

did you use the formula V = (1/3)*s^2*h ?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

what did you use for s and h?

OpenStudy (anonymous):

9 and 10

jimthompson5910 (jim_thompson5910):

V = (1/3)*s^2*h V = (1/3)*9^2*10 V = ??

OpenStudy (anonymous):

oops this time i got 27

jimthompson5910 (jim_thompson5910):

close

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i got 27.00

jimthompson5910 (jim_thompson5910):

that's incorrect, but close

OpenStudy (anonymous):

i did 81 times 10 times 1/3

OpenStudy (anonymous):

now i got 270 is that right

jimthompson5910 (jim_thompson5910):

pyramid A: 270 pyramid B: 1350

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i got 5.00 when i divided them

OpenStudy (anonymous):

am i supposed to do that?

jimthompson5910 (jim_thompson5910):

So B is 500% times bigger

jimthompson5910 (jim_thompson5910):

actually, hmm let me think for a second

OpenStudy (anonymous):

? wouldn't it be 5%

OpenStudy (anonymous):

or is it 20%

OpenStudy (anonymous):

@jim_thompson5910 u there

jimthompson5910 (jim_thompson5910):

are you given a list of choices?

OpenStudy (anonymous):

nope

jimthompson5910 (jim_thompson5910):

well since it's 5 times bigger, and 5 = 500%, that's why I'm thinking 500%

OpenStudy (anonymous):

thx